# InternScenes: A Large-scale Simulatable Indoor Scene Dataset with Realistic Layouts

### v1

@@ Abstract
The advancement of Embodied AI heavily relies on large-scale, simulatable 3D scene datasets characterized by scene diversity and realistic layouts. However, existing datasets typically suffer from limitations in data scale or diversity, sanitized layouts lacking small items, and severe object collisions. To address these shortcomings, we introduce InternScenes, a novel large-scale simulatable indoor scene dataset comprising approximately 40,000 diverse scenes by integrating three disparate scene sources, i.e., real-world scans, procedurally generated scenes, and designer-created scenes, including 1.96M 3D objects and covering 15 common scene types and 288 object classes. We particularly preserve massive small items in the scenes, resulting in realistic and complex layouts with an average of 41.5 objects per region. Our comprehensive data processing pipeline ensures simulatability by creating real-to-sim replicas for real-world scans, enhances interactivity by incorporating interactive objects into these scenes, and resolves object collisions by physical simulations. We demonstrate the value of InternScenes with two benchmark applications: scene layout generation and point-goal navigation. Both show the new challenges posed by the complex and realistic layouts. More importantly, InternScenes paves the way for scaling up the model training for both tasks, making the generation and navigation in such complex scenes possible. We commit to open-sourcing the data, models, and benchmarks to benefit the whole community.
 Github  |   data  |   Homepage
@@ 1 Introduction
Figure 1: InternScenes is a large-scale simulatable indoor scene dataset with diverse layout and various 3D objects. It suports tasks like layout generation and vision navigation.
@@
Consequently, as shown in Figure 1 our dataset consists of approximately $40,000$ diverse indoor scenes, including 48k regions from 15 common types in daily life, e.g., living region, resting region, dining region, etc., and features 1.96M objects and 800k CAD models covering a comprehensive taxonomy of 288 object classes within indoor scenes. Furthermore, we substitute roughly $20\%$ 3D assets inside with interactive objects from PartNet-Mobility and subsequently put all the scenes into the physical simulator to prevent object collisions, yielding a large-scale dataset of simulatable scenes with complex and realistic layouts. For example, each region of InternScenes has the highest-ever average number of 41.5 objects.
2 Related Work
Real-world Scans of Indoor 3D Scenes. To directly obtain information from the 3D world for perception, researchers have employed various sensors to scan real environments, capturing RGB-D images that are subsequently reconstructed and annotated. Datasets such as ScanNet, MP3D, and 3RScan (Dai et al., 2017; Wald et al., 2019; Chang et al., 2017; Yeshwanth et al., 2023; Straub et al., 2019; Ramakrishnan et al., 2021) are utilized to enhance models’ 3D perception capabilities. Although this direct scanning method preserves much of the real-world information, it is limited by the constraints of the collection equipment and the complexity of the data acquisition process. This often introduces noise, which challenges the accuracy of scene reconstruction and annotation. In recent years, researchers have made significant progress in enhancing reconstruction quality and annotation precision. For example, ScanNet++ (Yeshwanth et al., 2023) utilizes higher precision equipment compared to ScanNet to achieve improved reconstruction and semantic annotation. EmbodiedScan (Wang et al., 2024) has enriched datasets from ScanNet, MP3D, and 3RScan with extensive annotation information, including annotations for small objects within scenes. It expands the object categories to 288 and provides annotations with 9DoF bounding box information. Despite employing higher precision collection equipment and more detailed annotations, there is still a considerable gap between these scenes in simulation environments and real-world scenarios, along with inevitable annotation errors. Constructing a large number of finely annotated real-to-sim scenes is labor-intensive and time-consuming. As a result, researchers are increasingly focusing on indoor simulation scenes.
Simulated Indoor Scenes. To efficiently and cost-effectively obtain large-scale, detailed indoor scene data, researchers increasingly rely on computer software to construct and process synthetic indoor scenes (Li et al., 2018; Zheng et al., 2020; Roberts et al., 2021; Fu et al., 2021; Li et al., 2023; Deitke et al., 2022; Avetisyan et al., 2024), enabling the acquisition of multimodal data from diverse viewpoints. However, due to copyright restrictions, researchers often cannot access the original 3D assets directly and must rely on pre-rendered datasets. Although Hypersim (Roberts et al., 2021) provides a processing pipeline that allows users to purchase the original assets and follow the whole pipeline for custom rendering, this approach is prohibitively expensive, with costs reaching approximately \$57K. As an alternative, rule-based generative models such as SceneScript (Avetisyan et al., 2024) and Infinigen (Raistrick et al., 2024) can automatically generate an unlimited amount of 3D scene data via scripting. However, they are computationally intensive and time-consuming, and the resulting scenes often suffer from limited diversity. Another distinct approach is taken by 3D-FRONT (Fu et al., 2021), which has released 18K curated scene layouts and 13K CAD models, allowing researchers to reconstruct complete indoor scenes for novel tasks. However, since these layout designs are generated by learning from professional designers’ inspirations, they often lack realism and diversity, resulting in scenes with fewer objects and missing many small items that are common in real environments. In contrast, InternScenes comprises approximately 40K diverse indoor scenes, encompassing 48K regions across 15 common daily-life categories. Each region contains an average of 41.5 objects, indicating a high density of objects within our scenes, including small items. Moreover, around $30\%$ of the objects in each region are interactive.
@@ 2 Related Work
Real-to-Sim 3D Scene Generation. To construct scene datasets with authentic spatial distributions suitable for physics-based simulation and embodied AI training, researchers typically employ a real-to-sim paradigm to assemble scene datasets. In this approach, real environments are scanned to acquire detailed layout information, which is subsequently transformed into synthetic scene assets. For instance, the OpenRooms (Li et al., 2021) dataset builds upon ScanNet (Dai et al., 2017) indoor point-cloud, it aligns ShapeNet (Chang et al., 2015) CAD models with the scanned furniture by the Scan2CAD (Avetisyan et al., 2019) method. During this process, each object’s bounding box is meticulously refined to enforce orthogonality with both the floor and wall planes and to remove any floating or intersecting artifacts, resulting in physically coherent scene layouts and object placements. In contrast, certain methods forego point-cloud acquisition entirely, instead inferring spatial priors directly from a single image to synthesize 3D scenes. MIDI (Huang et al., 2024) conditions on a single image to generate multiple object assets in one pass. However, it frequently introduces visible artifacts and suffers from severe entanglement among objects of disparate scales, undermining realistic interactions. ACDC (Dai et al., 2024) method leverages vision-language models to extract scene distributions from a single image and reconstruct environments using BEHAVIOR-1K (Li et al., 2023) assets. Despite its promise, ACDC struggles to accurately represent complex scenes populated with numerous small objects, limiting its fidelity in such scenarios.
@@ 3.1 Multi-Source Data Processing
InternScenes-Real2Sim: Real-to-Sim Replica Creation on Real-world Scanned Scenes. Currently, retrieving scenes from real to simulation environments still faces two core challenges. First, the layouts in real-world scenes exhibit high diversity and complexity, as residents often have personalized preferences for object arrangements within scenes. Second, real scenes commonly contain a large number of small objects, which are more heterogeneous in category, greater in quantity, and display significantly varied poses compared to large furniture items. To address these challenges, we propose an effective retrieval pipeline, illustrated in Figure 2. The detailed annotations of regions and numerous small objects in EmbodiedScan precisely meet our requirements, making it a valuable data source. To ensure sufficient object density within each region, we defined a set of rules to merge small, semantically similar regions, guaranteeing that each resulting region contains at least eight objects.
@@
Specifically, at the Regions level, a single house typically contains multiple functional regions, yet these regions are not clearly delineated. For example, in an apartment suite, the living region and cooking region might coexist in the same scene, but designers often fail to distinctly define their boundaries, rendering it impossible to implement automatic segmentation using standard algorithms. At the Instances level, there are both furniture sets that are physically combined in the scene and parts that theoretically should be combined but are not. For instance, a sofa and the pillows or magazines placed on it are defined by the designer as a single instance, while the table legs and tabletop are split into separate instances. This approach to organization not only leads to significant ambiguity in semantic instance judgment, but also results in potentially inaccurate bounding box dimensions. To address the aforementioned issues, we refined the definition of region types within the scenes and performed splitting or merging operations on instances to capture the finest layout distribution within each region. The overall processing pipeline is illustrated in Figure 3.
@@
Instance Annotation. To address the hierarchical disorder among objects in the original scenes, we relied on human judgment to determine whether objects needed splitting or merging. For this purpose, we built an instance annotation tool that allows annotators to freely navigate the 3D scene and locate target objects for evaluation and labeling. Based on these annotations, we wrote scripts to automatically perform object splitting and merging. Subsequently, we rendered the processed objects from six different viewpoints and feed these images into the InternVL to generate semantic labels automatically, which are then verified by human annotators. Based on the region and instance annotation results, we further extract the object coordinates, bounding box dimensions, and rotation Euler angles within each region, thereby forming the necessary layout information.
@@ 3.2 Physics-Aware Scene Composition
To prevent collision and clipping issues between objects in the scene, we perform physical simulation optimization on the scenes gotten in the previous step. Specifically, we first conduct fine-tuning of the bounding boxes of the objects and then place them into a simulator to perform final computational adjustments to achieve the final scene layout.
@@
Simulator Processing. To further enhance the physical plausibility of smaller objects in the region and avoid common issues such as object collision and floating, we import the optimized furniture layout into the SAPIEN (Xiang et al., 2020) engine for detailed physics simulation, after completing the bounding-box-based optimization for large furniture items. Specific implementation details regarding bounding box optimization and physics simulation can be found in the supplementary material.
@@ 4 Dataset Statistics
Layout and Objects Statistics. Our dataset comprises three subsets, totaling 39870 scenes and 48381 regions across 15 categories. Specifically, the InternScenes-Real2Sim subset contains 9833 regions, InternScenes-Gen contains 11454 regions, and InternScenes-Synthetic contains 27094 regions. In total, 1.96M objects from 288 categories are placed across all regions, sampled from our asset library of 80M CAD models. These objects are sampled from our asset library containing 80 million CAD models. On average, each region contains 41.5 objects.
@@
Region Statistics. Figure 7 (a) shows the distribution of 15 region types. Figure 7 (b) shows the distribution of region area (in m2).
Object Statistics. Figure 8 shows the distribution of objects across 288 categories. The five most frequent object categories are chair, toy, book, light, and bottle. We also conduct a statistical study of the volume distribution of all object bounding boxes in the scenes (Figure 9(a)), and further analyzed the volume distributions of five representative object categories. These categories were selected to represent objects of varying scales, ranging from large furniture to small items: chair, bed, couch, bottle, and book (Figure 9(b)).
Figure 7: Region statistics.
Figure 8: Distribution of objects across 288 categories
Figure 9: Object bounding boxes volume statistics
@@
Region and Object Joint Statistics. Figure 10 illustrates the distribution of object density, measured as the number of objects per square meter (m2), across various regions. The average object density computed across all scenes is 1.296 objects per square meter.
In Figure 11, we show the distribution of 100 object categories across 15 regions, where the depth of the rectangle’s color and its size are positively correlated with the quantity of that object category within the corresponding region. The darker and larger the rectangle, the higher the frequency of that object category in the area.
@@ 5.1 Interior Scene Generation
Table 2: Quantitative evaluation results of ATISS, Diffuscene, and Physcene trained separately on the full and simplified versions of the Internscenes dataset. For SCA, the score closer to $50\%$ is better. Lower FID and CKL demonstrate better generation performance.
@@
ATISS(Paschalidou et al., 2021) 101.85 95.65 0.178 104.48 96.95 0.091 133.20 99.44 0.151
Diffuscene(Tang et al., 2024) 96.56 95.40 0.232 107.49 96.66 0.149 122.95 97.54 0.235
Physcene(Yang et al., 2024) 88.02 94.97 0.175 66.59 96.45 0.123 130.39 98.91 0.081
@@
ATISS(Paschalidou et al., 2021) 23.20 59.80 0.133 30.49 70.95 0.056 30.89 64.72 0.063
Diffuscene(Tang et al., 2024) 22.88 57.70 0.117 23.54 64.30 0.057 28.70 59.99 0.095
Physcene(Yang et al., 2024) 23.78 68.45 0.142 24.75 64.40 0.058 26.76 66.82 0.047
@@ 5.2 Navigation
Next, we choose point-goal navigation as the benchmark application of InternScenes for embodied AI. Previous scene datasets for point-goal navigation either has simple layout complexity or limited diversity. In contrast, InternScenes provides diverse simulation-ready environments that can generate considerable episodes therein. More importantly, it offers a challenging testbed for testing point-goal navigation algorithms in diverse, realistic, cluttered scenes.
@@
Experiment Setup. To evaluate the efficacy of our scene datasets for downstream Embodied AI tasks, we build a physically and visually realistic point-goal navigation benchmark based on IsaacSim and our scene assets, which distinguishes from prior physical-agnostic navigation benchmark, such as Habitat-Sim (Savva et al., 2019) and AI2Thor (Kolve et al., 2017). For a more comprehensive study on evaluating the sim-to-real gap of navigation approaches, we manually select 20 InternScenes-Real2Sim, 10 InternScenes-Gen based on layout complexity and quality. The wheeled robot ClearPath Dingo is considered as the navigation agent. Two metrics are considered in the benchmark: Success Rate and SPL. The former evaluates whether the agent can successfully find the a valid path leading to the goal. The latter evaluates the efficiency of the executed path compared with the oracle shortest path. Each scene is evaluated for 20 episodes, and we records an average distance between all starting points and target points to indicate the task difficulty.
@@
Baseline. Three representative baseline methods are considered in the evaluation. The first is an RL-based method DD-PPO (Wijmans et al., ) which is massively trained in Habitat-Sim (Savva et al., 2019). As DD-PPO trains the policy with respect to discrete action space, we deploy it in the continuous action space by multiply the discrete predict coordinates with a coefficient into linear and angular speed. The second is a pretrained diffusion-based imitation learning method NavDP (Cai et al., 2025). The third is a fine-tuned version of NavDP. To fine-tune the NavDP, we follow their data generation pipeline with our Internscenes assets, and composing a new navigation dataset with 118,784 trajectories.
@@
Results and Analysis. The navigation performance metrics are reported in the Table 3. The DD-PPO (Wijmans et al., ) achieves low success rate across all scenes, this implies the RL-based policy owns limited generalization ability when facing the continuous world and domain-gap in motion process. Although NavDP policy can select a best trajectory based on a pretrained critic function and accomplish navigation tasks in many scenarios, but the cluttered layout in Internscenes proposed unique challenges, thus results in $\sim$50% in success rate. By feeding the NavDP with additional navigation trajectories within Internscene, NavDP can slightly improve the overall performance. This implies the diversity of our Internscene dataset can benefit the model training, but how to scale up the model’s capacity with increasing datasets is still an unsolved problem.
@@
Discussion and Conclusion. Based on the navigation performance metrics, we find significant performance desendent in our evaluation framework. By diving into the failure cases, we conclude three main challenges in our benchmark and proposes potential direction for future navigation method research. Firstly, our realistic scene assets tends to exhibit cluttered layouts in rooms, which requires more accurate path planning ability and recovery ability from collision. The absense of failure recovery ability in the baseline methods is an important reason causing the performance drop in cluttered environments. Secondly, our scene assets often demonstrates narrow pathways where the accessibility is totally depend on the robot embodiment information. But as most learning-based navigation methods only depend on the exteroception observations, this limits the navigation performance. Thirdly, real-world objects often own connected parts that are tiny but regarded as obstacles, such as legs of office chairs. Such tiny obstacles may be captured in the visual observations in limited frames, but are essential for safe path planning. And this proposes great challenge for spatial perception ability of the navigation approaches. The aforementioned three features makes our navigation benchmark becomes an ideal platform to help evaluate the sim-to-real gap of navigation approaches. More qualitative results and visualization are provided in the supplementary mateirals.

### v2

@@ Abstract
The advancement of Embodied AI heavily relies on large-scale, simulatable 3D scene datasets characterized by scene diversity and realistic layouts. However, existing datasets typically suffer from limitations in data scale or diversity, sanitized layouts lacking small items, and severe object collisions. To address these shortcomings, we introduce InternScenes, a novel large-scale simulatable indoor scene dataset comprising approximately 40,000 diverse scenes by integrating three disparate scene sources, i.e., real-world scans, procedurally generated scenes, and designer-created scenes, including 1.96M 3D objects and covering 15 common scene types and 288 object classes. We particularly preserve massive small items in the scenes, resulting in realistic and complex layouts with an average of 41.5 objects per region. Our comprehensive data processing pipeline ensures simulatability by creating real-to-sim replicas for real-world scans, enhances interactivity by incorporating interactive objects into these scenes, and resolves object collisions by physical simulations. We demonstrate the value of InternScenes with two benchmark applications: scene layout generation and point-goal navigation. Both show the new challenges posed by the complex and realistic layouts. More importantly, InternScenes paves the way for scaling up the model training for both tasks, making the generation and navigation in such complex scenes possible. We commit to open-sourcing the data and benchmarks to benefit the whole community.
Code | Data | Homepage
Figure 1: InternScenes is a large-scale, simulatable indoor scene dataset with diverse layouts and various 3D objects. It supports various tasks, such as scene layout generation and vision navigation.
@@ 1 Introduction
Consequently, as shown in Figure 1 our dataset consists of approximately $40,000$ diverse indoor scenes, including 48k regions from 15 common types in daily life, e.g., living region, resting region, dining region, etc., and features 1.96M objects and 800k CAD models covering a comprehensive taxonomy of 288 object classes within indoor scenes. Furthermore, we substitute roughly $20\%$ 3D assets inside with interactive objects from PartNet-Mobility and subsequently put all the scenes into the physical simulator to prevent object collisions, yielding a large-scale dataset of simulatable scenes with complex and realistic layouts. For example, each region of InternScenes has the highest-ever average number of 41.5 objects. Table. 1 provides an overall comparison between InternScenes and existing 3D indoor scene datasets.
2 Related Work
Real-world Scans of Indoor 3D Scenes. To directly obtain information from the 3D world for perception, researchers have employed various sensors to scan real environments, capturing RGB-D images that are subsequently reconstructed and annotated. Datasets (Dai et al., 2017; Wald et al., 2019; Chang et al., 2017; Yeshwanth et al., 2023; Straub et al., 2019; Ramakrishnan et al., 2021) such as ScanNet, MP3D, and 3RScan are utilized to enhance models’ 3D perception capabilities. Although this direct scanning method preserves much of the real-world information, it is limited by the constraints of the collection equipment and the complexity of the data acquisition process. This often introduces noise, which challenges the accuracy of scene reconstruction and annotation.
In recent years, researchers have made significant progress in enhancing reconstruction quality and annotation precision. For example, ScanNet++ (Yeshwanth et al., 2023) utilizes higher precision equipment compared to ScanNet to achieve improved reconstruction and semantic annotation. EmbodiedScan (Wang et al., 2024) has enriched datasets from ScanNet, MP3D, and 3RScan with extensive annotation information, including annotations for small objects within scenes. It expands the object categories to 288 and provides annotations with 9DoF bounding box information. Despite employing higher precision collection equipment and more detailed annotations, there is still a considerable gap between these scenes in simulation environments and real-world scenarios, along with inevitable annotation errors. Constructing a large number of finely annotated real-to-sim scenes is labor-intensive and time-consuming. As a result, researchers are increasingly focusing on indoor simulation scenes.
Simulated Indoor Scenes. To efficiently and cost-effectively obtain large-scale, detailed indoor scene data, researchers increasingly rely on computer software to construct and process synthetic indoor scenes (Li et al., 2018; Zheng et al., 2020; Roberts et al., 2021; Fu et al., 2021; Li et al., 2023; Deitke et al., 2022; Avetisyan et al., 2024), enabling the acquisition of multimodal data from diverse viewpoints. However, due to copyright restrictions, researchers often cannot access the original 3D assets directly and must rely on pre-rendered datasets. Although Hypersim (Roberts et al., 2021) provides a processing pipeline that allows users to purchase the original assets and follow the whole pipeline for custom rendering, this approach is prohibitively expensive, with costs reaching approximately \$57K.
As an alternative, rule-based generative models such as SceneScript (Avetisyan et al., 2024) and Infinigen (Raistrick et al., 2024) can automatically generate an unlimited amount of 3D scene data via scripting. However, they are computationally intensive and time-consuming, and the resulting scenes often suffer from limited diversity. Another distinct approach is taken by 3D-FRONT (Fu et al., 2021), which has released 18K curated scene layouts and 13K CAD models, allowing researchers to reconstruct complete indoor scenes for novel tasks. However, since these layout designs are generated by learning from professional designers’ inspirations, they often lack realism and diversity, resulting in scenes with fewer objects and missing many small items that are common in real environments. In contrast, InternScenes comprises approximately 40K diverse indoor scenes, encompassing 48K regions across 15 common daily-life categories. Each region contains an average of 41.5 objects, indicating a high density of objects within our scenes, including small items. Moreover, around $20\%$ of the objects in each region are interactive.
@@ 2 Related Work
Real-to-Sim 3D Scene Generation. To construct scene datasets with authentic spatial distributions suitable for physics-based simulation and embodied AI training, researchers typically employ a real-to-sim paradigm to assemble scene datasets. In this approach, real environments are scanned to acquire detailed layout information, which is subsequently transformed into synthetic scene assets. For instance, the OpenRooms (Li et al., 2021) dataset builds upon ScanNet (Dai et al., 2017) indoor point cloud; it aligns ShapeNet (Chang et al., 2015) CAD models with the scanned furniture by the Scan2CAD (Avetisyan et al., 2019) method. During this process, each object’s bounding box is meticulously refined to enforce orthogonality with both the floor and wall planes and to remove any floating or intersecting artifacts, resulting in physically coherent scene layouts and object placements. Concurrent work MetaScenes (Yu et al., 2025) also builds delicate Real2Sim replica via replacing objects in ScanNet (Dai et al., 2017) Scenes, but its coverage is limited in a single data source. In contrast, certain methods forego point‐cloud acquisition entirely, instead inferring spatial priors directly from a single image to synthesize 3D scenes. MIDI (Huang et al., 2024) conditions on a single image to generate multiple object assets in one pass. However, it frequently introduces visible artifacts and suffers from severe entanglement among objects of disparate scales, undermining realistic interactions. ACDC (Dai et al., 2024) method leverages vision-language models to extract scene distributions from a single image and reconstruct environments using BEHAVIOR-1K (Li et al., 2023) assets. Despite its promise, ACDC struggles to accurately represent complex scenes populated with numerous small objects, limiting its fidelity in such scenarios.
@@ 3.1 Multi-Source Data Processing
InternScenes-Real2Sim: Real-to-Sim Replica Creation on Real-world Scanned Scenes. Currently, retrieving scenes from real to simulation environments still faces two core challenges. First, the layouts in real-world scenes exhibit high diversity and complexity, as residents often have personalized preferences for object arrangements within scenes. Second, real scenes commonly contain a large number of small objects, which are more heterogeneous in category, greater in quantity, and display significantly varied poses compared to large furniture items. To address these challenges, we propose an effective retrieval pipeline, illustrated in Figure 2. The detailed annotations of regions and numerous small objects in EmbodiedScan precisely meet our requirements, making it a valuable data source. To ensure sufficient object density within each region, we defined a set of rules to merge small, semantically similar regions, guaranteeing that each resulting region contains at least $8$ objects.
@@
Specifically, at the Regions level, a single house typically contains multiple functional regions, yet these regions are not clearly delineated. For example, in an apartment suite, the living region and cooking region might coexist in the same scene, but designers often fail to distinctly define their boundaries, rendering it impossible to implement automatic segmentation using standard algorithms. At the Instances level, there are both furniture sets that are physically combined in the scene and parts that theoretically should be combined but are not. For instance, a sofa and the pillows or magazines placed on it are defined by the designer as a single instance, while the table legs and tabletop are split into separate instances. This approach to organization not only leads to significant ambiguity in semantic instance judgment but also results in potentially inaccurate bounding box dimensions. To address the aforementioned issues, we refined the definition of region types within the scenes and performed splitting or merging operations on instances to capture the finest layout distribution within each region. The overall processing pipeline is illustrated in Figure 3.
@@
Instance Annotation. To address the hierarchical disorder among objects in the original scenes, we relied on human judgment to determine whether objects needed splitting or merging. For this purpose, we built an instance annotation tool that allows annotators to freely navigate the 3D scene and locate target objects for evaluation and labeling. Based on these annotations, we wrote scripts to automatically perform object splitting and merging. Subsequently, we rendered the processed objects from six different viewpoints and fed these images into the InternVL to generate semantic labels automatically, which are then verified by human annotators. Based on the region and instance annotation results, we further extract the object coordinates, bounding box dimensions, and rotation Euler angles within each region, thereby forming the necessary layout information.
@@ 3.2 Physics-Aware Scene Composition
To prevent collision and clipping issues between objects in the scene, we perform physical simulation optimization on the scenes obtained in the previous step. Specifically, we first conduct fine-tuning of the bounding boxes of the objects and then place them into a simulator to perform final computational adjustments to achieve the final scene layout.
@@
Simulator Processing. To further enhance the physical plausibility of smaller objects in the region and avoid common issues such as object collision and floating, we import the optimized furniture layout into the SAPIEN (Xiang et al., 2020) engine for detailed physics simulation, after completing the bounding-box-based optimization for large furniture items. Implementation details regarding bounding box optimization and physics simulation can be found in the supplementary material.
@@ 4 Dataset Statistics
Figure 7: Region statistics. Our dataset includes 15 common scene categories, such as the resting room and the living room. We also show the distribution of region areas.
Figure 8: Distribution of objects across 288 categories. We list the 30 categories with the highest occurrence frequency.
Figure 9: Object bounding boxes volume statistics.
@@
Region and Objects Statistics. Our dataset comprises three subsets, totaling 39870 scenes and 48381 regions across 15 categories. Specifically, the InternScenes-Real2Sim subset contains 9833 regions, InternScenes-Gen contains 11454 regions, and InternScenes-Synthetic contains 27094 regions. In total, 1.96M objects from 288 categories are placed across all regions, sampled from our asset library of 80M CAD models. These objects are sampled from our asset library containing 80 million CAD models. On average, each region contains 41.5 objects.
Figure 7 (a) shows the distribution of 15 region types. Figure 7 (b) shows the distribution of region area (in m2). Figure 8 shows the distribution of objects across 288 categories. The five most frequent object categories are chair, toy, book, light, and bottle. We also conduct a statistical study of the volume distribution of all object bounding boxes in the scenes (Figure 9(a)), and further analyzed the volume distributions of five representative object categories. These categories were selected to represent objects of varying scales, ranging from large furniture to small items: chair, bed, couch, bottle, and book (Figure 9(b)).
Figure 10 illustrates the distribution of object density, measured as the number of objects per square meter (m2), across various regions. The average object density computed across all scenes is 1.296 objects per square meter. In Figure 11, we show the distribution of 100 object categories across 15 regions, where the depth of the rectangle’s color and its size are positively correlated with the quantity of that object category within the corresponding region. The darker and larger the rectangle, the higher the frequency of that object category in the area.
@@ 5.1 Interior Scene Generation
Table 2: Quantitative evaluation results of ATISS(Paschalidou et al., 2021), DiffuScene(Tang et al., 2024), and PhyScene(Yang et al., 2024) trained separately on the full and simplified versions of the Internscenes dataset. For SCA, the score closer to $50\%$ is better. Lower FID and CKL demonstrate better generation performance.
@@
ATISS 101.85 95.65 0.178 104.48 96.95 0.091 133.20 99.44 0.151
DiffuScene 96.56 95.40 0.232 107.49 96.66 0.149 122.95 97.54 0.235
PhyScene 88.02 94.97 0.175 66.59 96.45 0.123 130.39 98.91 0.081
@@
ATISS 23.20 59.80 0.133 30.49 70.95 0.056 30.89 64.72 0.063
DiffuScene 22.88 57.70 0.117 23.54 64.30 0.057 28.70 59.99 0.095
PhyScene 23.78 68.45 0.142 24.75 64.40 0.058 26.76 66.82 0.047
@@ 5.2 Navigation
Next, we choose point-goal navigation as the benchmark application of InternScenes for embodied AI. Previous scene datasets for point-goal navigation either have simple layout complexity or limited diversity. In contrast, InternScenes provides diverse simulation-ready environments that can generate considerable episodes therein. More importantly, it offers a challenging testbed for testing point-goal navigation algorithms in diverse, realistic, cluttered scenes.
@@
Experiment Setup. To evaluate the efficacy of our scene datasets for downstream Embodied AI tasks, we build a physically and visually realistic point-goal navigation benchmark based on IsaacSim and our scene assets, which distinguishes from prior physical-agnostic navigation benchmarks, such as Habitat-Sim (Savva et al., 2019) and AI2Thor (Kolve et al., 2017). For a more comprehensive investigation of the sim-to-real gap in navigation approaches, we manually select 20 scenes from InternScenes-Real2Sim and 10 from InternScenes-Gen, with selection criteria based on layout complexity and asset quality. The wheeled robot ClearPath Dingo is adopted as the navigation agent. Two metrics are employed in the benchmark: Success Rate and Shortest Path Length (SPL). Success Rate assesses whether the agent can find a valid path to reach the goal, while SPL measures the efficiency of the executed path relative to the oracle shortest path. Each scene is evaluated across 20 episodes, and we report the average distance between all starting points and target points to quantify task difficulty.
@@
Baseline. Three representative baseline methods are considered in the evaluation. The first is an RL-based method, DD-PPO (Wijmans et al., ), which is massively trained in Habitat-Sim (Savva et al., 2019). As DD-PPO trains the policy with respect to a discrete action space, we deploy it in the continuous action space by multiplying the discrete predicted coordinates with a coefficient into linear and angular speed. The second is a pretrained diffusion-based imitation learning method, NavDP (Cai et al., 2025). The third is a fine-tuned version of NavDP. To fine-tune the NavDP, we follow their data generation pipeline with our Internscenes assets and compose a new navigation dataset with 118,784 trajectories.
@@
Results and Analysis. Navigation performance metrics are reported in Table 3. DD-PPO (Wijmans et al., ) achieves a low success rate across all scenes, indicating that RL-based policies have limited generalization capabilities when confronted with continuous action spaces and domain gaps during motion execution. While NavDP (Cai et al., 2025) can select optimal trajectories using its pretrained critic function and complete navigation tasks in many scenarios, the cluttered layouts in InternScenes pose unique challenges, leading to a success rate of approximately $50\%$. By fine-tuning NavDP with additional navigation trajectories from InternScenes, we observe a slight improvement in its overall performance. This result demonstrates that the diversity of our InternScenes dataset can facilitate model training; however, how to scale model capacity to leverage increasingly large datasets remains an open problem.
@@
Discussion and Conclusion. Based on the navigation performance metrics, we observe a significant performance decline in our evaluation framework. By analyzing failure cases in depth, we identify three key challenges in our benchmark and propose a potential direction for future research on navigation methods. First, the realistic scene assets in our benchmark tend to feature cluttered room layouts, which demand more precise path-planning capabilities and collision recovery mechanisms. The lack of collision recovery capabilities in the baseline methods is a key factor contributing to their performance degradation in cluttered environments. Second, our scene assets frequently include narrow pathways, where traversability depends entirely on the robot’s embodiment information. However, most learning-based navigation methods rely solely on exteroceptive observations, which constrains their navigation performance in such scenarios. Third, real-world objects often have small connected components (e.g., office chair legs) that are classified as obstacles. While these tiny obstacles may be captured in visual observations only in limited frames, they are critical for safe path planning—posing a major challenge to the spatial perception capabilities of navigation approaches. These three characteristics make our navigation benchmark an ideal platform for evaluating the sim-to-real gap of navigation methods.